what is the equation of a circle passing through (6,2) and tangent to the line x-4y-15=0 at (3,-3)

1
Expert's answer
2021-12-16T16:11:55-0500

The equation of the circle with center "(a, b)" and radius "r" is



"(x-a)^2+(y-b)^2=r^2"

Differentiate both sides with respect to "x"



"2(x-a)+2(y-b)y'=0"

Solve for "y'"



"y'=-\dfrac{x-a}{y-b}"

The slope of the line tangent to the circle at "(3, -3)"



"slope=m=-\dfrac{3-a}{-3-b}=\dfrac{3-a}{3+b}"

The equation of the tangent line tangent to the circle at "(3, -3)" is



"y-(-3)=\dfrac{3-a}{3+b}(x-3)""y=\dfrac{3-a}{3+b}x-3\cdot\dfrac{3-a}{3+b}-3""y=\dfrac{3-a}{3+b}x-3\cdot\dfrac{6+b-a}{3+b}"

The line "x-4y-15=0"



"y=\dfrac{1}{4}x-\dfrac{15}{4}"

Then



"\begin{cases}
 \dfrac{3-a}{3+b}=\dfrac{1}{4}\\
\\
 -3\cdot\dfrac{6+b-a}{3+b}=-\dfrac{15}{4}
\end{cases}""\begin{cases}
b=9-4a\\
\\
 \dfrac{6+b-a}{3-a}=5
\end{cases}""b=9-4a, a\in \R"


Substitute



"(x-a)^2+(y-9+4a)^2=r^2"


Point "(3,-3)"



"(3-a)^2+(-3-9+4a)^2=r^2""9-6a+a^2+144-96a+16a^2=r^2""r^2=17a^2-102a+153"

Point "(6,2)"



"(6-a)^2+(2-9+4a)^2=r^2""36-12a+a^2+49-56a+16a^2=r^2""r^2=17a^2-68a+85""r^2=17a^2-102a+153=17a^2-68a+85""34a=68""a=2""b=9-4(2)=1""r^2=17(2)^2-68(2)+85=17"

The equation of the circle is



"(x-2)^2+(y-1)^2=17"